# Congestion-Approximators from the Bottom Up

### v2

@@ 4 Bottom-Up Congestion-Approximator
which satisfies $\mathbf{x}^{\prime}\leq\deg_{\partial\mathcal{R}_{\geq i+2}}$. By induction, there exists a vector $\mathbf{y}^{\prime}\in\mathbb{R}^{V}_{\geq 0}$ with $\mathbf{y}^{\prime}\leq\deg_{\partial\mathcal{P}_{i+1}}$ and a flow $f_{1}$ routing demand $\mathbf{x}^{\prime}-\mathbf{y}^{\prime}$ with congestion $(2L-1-2(i+1))\beta$.
Next, consider the vector $\mathbf{x}-\mathbf{x}^{\prime}+\mathbf{y}^{\prime}$, which satisfies $\mathbf{x}-\mathbf{x}^{\prime}+\mathbf{y}^{\prime}\geq\mathbf{x}-\mathbf{x}^{\prime}\geq\mathbf{0}$, where the second inequality holds by Claim 4.3. If $\deg_{\partial\mathcal{R}_{\geq i+1}}(v)=0$, then $\mathbf{x}(v)=0$ and $\mathbf{x}^{\prime}(v)=0$, so $(\mathbf{x}-\mathbf{x}^{\prime}+\mathbf{y}^{\prime})(v)=\mathbf{y}^{\prime}(v)\leq\deg_{\partial\mathcal{P}_{i+1}}(v)$. Otherwise, if $\deg_{\partial\mathcal{R}_{\geq i+1}}(v)>0$, then
$(\mathbf{x}-\mathbf{x}^{\prime}+\mathbf{y}^{\prime})(v)$ $=\bigg{(}\frac{\deg_{\partial\mathcal{R}_{\geq i+1}}(v)-\deg_{\partial\mathcal{R}_{\geq i+2}}(v)}{\deg_{\partial\mathcal{R}_{\geq i+1}}(v)}\bigg{)}\mathbf{x}(v)+\mathbf{y}^{\prime}(v)$
$\leq\frac{\deg_{\partial\mathcal{P}_{i+1}}(v)}{\deg_{\partial\mathcal{R}_{\geq i+1}}(v)}\mathbf{x}(v)+\mathbf{y}^{\prime}(v)$
$\leq\deg_{\partial\mathcal{P}_{i+1}}(v)+\deg_{\partial\mathcal{P}_{i+1}}(v),$
where the first inequality holds by Claim 4.4. Combining the two cases, we obtain $\mathbf{x}-\mathbf{x}^{\prime}+\mathbf{y}^{\prime}\leq 2\deg_{\partial\mathcal{P}_{i+1}}$. By Claim 4.8 on vector $(\mathbf{x}-\mathbf{x}^{\prime}+\mathbf{y}^{\prime})/2$, there exists a vector $\mathbf{t}\in\mathbb{R}^{V}_{\geq 0}$ with $\mathbf{t}\leq\mathrm{deg}_{\partial\mathcal{P}_{i}}/2$ and a flow routing demand $(\mathbf{x}-\mathbf{x}^{\prime}+\mathbf{y}^{\prime})/2-\mathbf{t}$ with congestion $\beta$. Scaling this flow by factor $2$, we obtain a flow $f_{2}$ routing demand $\mathbf{x}-\mathbf{x}^{\prime}+\mathbf{y}^{\prime}-2\mathbf{t}$ with congestion $2\beta$. The final flow is the sum of flows $f_{1}$ and $f_{2}$, which routes demand $(\mathbf{x}^{\prime}-\mathbf{y}^{\prime})+(\mathbf{x}-\mathbf{x}^{\prime}+\mathbf{y}^{\prime}-2\mathbf{t})=\mathbf{x}-2\mathbf{t}$ and has congestion $(2L-1-2(i+1))\beta+2\beta=(2L-1-2i)\beta$. We set $\mathbf{y}=2\mathbf{t}\leq\deg_{\partial\mathcal{P}_{i}}$, which concludes the inductive step and hence the proof. ∎
Claim 4.10.
@@
For any $i\in[L-1]$, consider any vector $\mathbf{x}\in\mathbb{R}^{V}_{\geq 0}$ with $\mathbf{x}\leq\deg_{\partial\mathcal{R}_{\geq i+1}}$. There exists a vector $\mathbf{y}\in\mathbb{R}^{V}_{\geq 0}$ such that
@@
    $\mathbf{y}\leq\textup{deg}_{\partial\mathcal{P}_{i}}+2L\beta\deg_{\partial\mathcal{P}_{i+1}}$.
@@
Apply Claim 4.9 on vector $\mathbf{x}$, obtaining a vector $\mathbf{y}\in\mathbb{R}^{V}_{\geq 0}$ which we relabel as $\mathbf{y}^{\prime}$, and a flow $f$ routing demand $\mathbf{x}-\mathbf{y}^{\prime}$ with congestion $(2L-1-2i)\beta\leq 2L\beta$. Take a path decomposition of flow $f$ where each vertex is the start of $\mathbf{x}(v)$ total capacity of (potentially empty) paths and the end of $\mathbf{y}^{\prime}(v)$ total capacity of (potentially empty) paths. For each path starting at a vertex $v$ in some cluster $C\in\mathcal{P}_{i+1}$, perform the following operation. If the path contains an edge $(u,v)$ in $\partial C$ with $u\in C$, then replace the path with its prefix ending at $u$; otherwise, do nothing to the path. Note that the modified path ends in the same cluster $C\in\mathcal{P}_{i+1}$ as its starting point. These modified paths form a new flow $f^{\prime}$, which also has congestion $2L\beta$.
@@
Define vector $\mathbf{y}\in\mathbb{R}^{V}_{\geq 0}$ such that each vertex $u\in V$ is the endpoint of $\mathbf{y}(u)$ total capacity of paths in the new flow $f^{\prime}$. In other words, the flow $f^{\prime}$ routes demand $\mathbf{x}-\mathbf{y}$. We have shown that $\mathbf{y}\leq\mathbf{y}^{\prime}+2L\beta\deg_{\partial\mathcal{P}_{i+1}}$, and combined with the guarantee $\mathbf{y}^{\prime}\leq\deg_{\partial\mathcal{P}_{i}}$ from Claim 4.9, we conclude that $\mathbf{y}\leq\deg_{\partial\mathcal{P}_{i}}+2L\beta\deg_{\partial\mathcal{P}_{i+1}}$. Finally, recall that each path of $f^{\prime}$ starts and ends in the same cluster of $\mathcal{P}_{i+1}$, so $(\mathbf{x}-\mathbf{y})(C)=0$ for all clusters $C\in\mathcal{P}_{i+1}$. ∎
@@
Since $|\mathbf{t}|\leq\deg_{\partial\mathcal{R}_{\geq i+1}}$, we apply Claim 4.10 with $\mathbf{x}=\mathbf{t}$, obtaining a vector $\mathbf{y}\in\mathbb{R}^{V}_{\geq 0}$ which we relabel as $\mathbf{t}^{\prime}$. We have $|\mathbf{t}^{\prime}|=\mathbf{t}^{\prime}\leq\deg_{\partial\mathcal{P}_{i}}+2L\beta\deg_{\partial\mathcal{P}_{i+1}}$ and $(\mathbf{t}-\mathbf{t}^{\prime})(C)=0$ for all clusters $C\in\mathcal{P}_{i+1}$. Let $f_{1}$ be the flow routing demand $\mathbf{t}-\mathbf{t}^{\prime}$ with congestion $2L\beta$.
Consider a cluster $C\in\mathcal{P}_{i+1}$. We have $(\mathbf{s}-\mathbf{t}^{\prime})(C)=(\mathbf{s}-\mathbf{t})(C)+(\mathbf{t}-\mathbf{t}^{\prime})(C)=0$. Moreover, for all vertices $v\in C$, we have
$|\mathbf{s}(v)-\mathbf{t}^{\prime}(v)|\leq|\mathbf{s}(v)|+|\mathbf{t}^{\prime}(v)|$ $\leq i\deg_{\partial\mathcal{P}_{i}}(v)+\deg_{\partial\mathcal{P}_{i}}(v)+2L\beta\deg_{\partial\mathcal{P}_{i+1}}(v)$
$=(i+1)\deg_{\partial\mathcal{P}_{i}}(v)+2L\beta\deg_{\partial C}(v)$
$\leq 3L\beta\deg_{\partial\mathcal{P}_{i}\cup\partial C}(v).$
We conclude that the scaled-down and restricted vector $\frac{1}{3L\beta}(\mathbf{s}-\mathbf{t}^{\prime})|_{C}$ is a demand satisfying $\big{|}\frac{1}{3L\beta}(\mathbf{s}-\mathbf{t}^{\prime})|_{C}\big{|}\leq\mathrm{deg}_{\partial\mathcal{P}_{i}\cup\partial C}|_{C}$. By assumption (2) of Lemma 4.2, the collection of vertex weightings $\{\mathrm{deg}_{\partial\mathcal{P}_{i}\cup\partial C}|_{C}\in\mathbb{R}^{V}_{\geq 0}:C\in\mathcal{P}_{i+1}\}$ mixes simultaneously in $G$ with congestion $\alpha$, so there is a flow in $G$ routing demand $\sum_{C\in\mathcal{P}_{i+1}}\frac{1}{3L\beta}(\mathbf{s}-\mathbf{t}^{\prime})|_{C}=\frac{1}{3L\beta}(\mathbf{s}-\mathbf{t}^{\prime})$ with congestion $\alpha$. Scaling this flow by factor $3L\beta$, we obtain a flow $f_{2}$ routing demand $\mathbf{s}-\mathbf{t}^{\prime}$ with congestion at most $3L\alpha\beta$.
The final flow $f$ is $f_{2}-f_{1}$, which routes demand $(\mathbf{s}-\mathbf{t}^{\prime})-(\mathbf{t}-\mathbf{t}^{\prime})=\mathbf{s}-\mathbf{t}$ and has congestion $2L\alpha+3L\alpha\beta\leq 5L\alpha\beta$, concluding the proof of Lemma 4.6. ∎
@@ 5 Partitioning Algorithm
The guarantee $\delta\mathcal{P}_{i+1}\leq\frac{1}{2}\delta\mathcal{P}_{i}$ ensures that for $L=O(\log(nW))$, we must have $\delta\mathcal{P}_{L}<1$. Since all edge capacities are assumed to be at least $1$, we conclude that $\delta\mathcal{P}_{L}=0$ and $\mathcal{P}_{L}=\{V\}$, fulfilling property (1) of Theorem 4.1. By Theorem 4.1, we obtain a congestion-approximator with quality $5L^{2}\alpha\beta=O(\log^{10}(nW))$. ∎
@@
At a high level, our algorithm proceeds similarly to expander decomposition algorithms like [34], where expanders are defined with respect to the vertex weighting $\deg_{\partial\mathcal{P}_{L}}$ of partition $\mathcal{P}_{L}$. We iteratively decompose the graph using the cut-matching game of [18]: for each cluster of the decomposition, we either compute a low-conductance cut or certify that the current cluster is mixing (or in [34] terms, a nearly expander). The matching step of the cut-matching game requires a call to approximate min-cut/max-flow, but recall that the partitions $\mathcal{P}_{1},\ldots,\mathcal{P}_{L}$ only form a pseudo-congestion-approximator. Luckily, we can modify the graph in a way that the pseudo-congestion-approximator can be adapted to an actual congestion-approximator for the new graph. We then black-box a cut/flow algorithm on this new graph, and then modify the cut and flow to fit the old graph in a way that suffices for the cut-matching game.
@@
    In Section 5.2, we cite the (approximate) fair cut/flow algorithm of [25], which computes a cut/flow pair with desirable properties given a congestion-approximator.
@@
    In Section 5.3, we introduce the cut-matching game as well as a trimming procedure similar to [34]. Both the cut-matching game and the trimming step use the fair cut/flow algorithm (Section 5.2) on the modified graphs for which we have a congestion-approximator (Section 5.1).
@@ 5.1 Congestion-approximator
For any demand $\mathbf{b}\in\mathbb{R}^{V}$ satisfying $|\mathbf{b}(C)|\leq\delta C$ for all $C\in\mathcal{C}$, there exists a demand $\mathbf{b}^{\prime}\in\mathbb{R}^{V}$ satisfying $|\mathbf{b}^{\prime}|\leq L\deg_{\partial\mathcal{P}_{L}}$ and a flow in $G$ routing demand $\mathbf{b}-\mathbf{b}^{\prime}$ with congestion $\kappa$, where we define $\kappa=5L^{2}\alpha\beta$.
Suppose we start with the entire graph $G$, and then add a vertex $x$ connected to each vertex $v\in V$ with an edge of capacity $\deg_{\partial\mathcal{P}_{L}}(v)$. In the setting of Fact 5.4, suppose we wish to route the demand $\mathbf{b}$. We start with the flow routing demand $\mathbf{b}-\mathbf{b}^{\prime}$ as promised by Fact 5.4. To route the remaining demand $\mathbf{b}^{\prime}$, we simply use the new edges incident to $x$: for each vertex $v\in V$, send $|\mathbf{b}^{\prime}(v)|\leq L\deg_{\partial\mathcal{P}_{L}}$ flow along the edge $(x,v)$ in the proper direction, which is a flow with congestion $L$. Overall, we obtain a flow routing demand $\mathbf{b}$ with congestion $\kappa+L$, and with some more work, we can show that $\mathcal{C}$ is a congestion approximator of the new graph.
@@
Let $\mathcal{C}|_{A}$ be the collection $\{C\cap A:C\in\mathcal{C}\}$. Then, $\mathcal{C}|_{A}\cup\{\{x\},\{s\},\{t\}\}$ is a congestion-approximator of $H$ with quality $O(\beta\gamma^{-1}(\kappa+L))$.
For the rest of Section 5.1, we prove Lemma 5.5. Consider a demand $\mathbf{b}\in\mathbb{R}^{A\cup\{x,s,t\}}$ satisfying $|\mathbf{b}(C\cap A)|\leq\delta_{H}(C\cap A)$ for all $C\in\mathcal{C}$ as well as $|\mathbf{b}(v)|\leq\delta_{H}\{v\}$ for $v\in\{x,s,t\}$. We want to establish a flow on $H$ routing demand $\mathbf{b}$ with congestion $O(\beta\gamma^{-1}(\kappa+L))$.
@@
so we can apply Fact 5.4 on demand $\widetilde{\mathbf{b}}^{\prime}/(1+2\beta+2\gamma)$ to obtain a demand $\mathbf{b}^{\prime}\in\mathbb{R}^{V}$ satisfying $|\mathbf{b}^{\prime}|\leq L\deg_{\partial_{G}\mathcal{P}_{L}}$ and a flow on $G$ routing demand $\widetilde{\mathbf{b}}^{\prime}/(1+2\beta+2\gamma)-\mathbf{b}^{\prime}$ with congestion $\kappa$. Scaling this flow by factor $(1+2\beta+2\gamma)$, we obtain a flow $f^{\prime}$ on $G$ routing demand $\widetilde{\mathbf{b}}^{\prime}-(1+2\beta+2\gamma)\mathbf{b}^{\prime}$ with congestion $(1+2\beta+2\gamma)\kappa$.
@@
Finally, since demand $(1+2\beta+2\gamma)\mathbf{b}^{\prime}$ satisfies $|(1+2\beta+2\gamma)\mathbf{b}^{\prime}|\leq(1+2\beta+2\gamma)L\deg_{\partial_{G}\mathcal{P}_{L}}$, we can directly route the demand along the edges $(v,x)$ of capacity $\gamma\deg_{\partial_{G}\mathcal{P}_{L}\cup\partial_{G}A}(v)$, which is a routing with congestion $\gamma^{-1}(1+2\beta+2\gamma)L$.
Adding up all three routings, we have routed the initial demand $\mathbf{b}$ with congestion $1+\gamma^{-1}(1+2\beta+2\gamma)\kappa+\gamma^{-1}(1+2\beta+2\gamma)L=O(\beta\gamma^{-1}(\kappa+L))$, concluding the proof.

### v3

@@ 4 Bottom-Up Congestion-Approximator
which satisfies $\mathbf{x}^{\prime}\leq\deg_{\partial\mathcal{R}_{\geq i+1}}$. By induction, there exists a vector $\mathbf{y}^{\prime}\in\mathbb{R}^{V}_{\geq 0}$ with $\mathbf{y}^{\prime}\leq 2\deg_{\partial\mathcal{P}_{i+1}}$ and a flow $f_{1}$ routing demand $\mathbf{x}^{\prime}-\mathbf{y}^{\prime}$ with congestion $(2L-2(i+1))\beta$. By Subclaim 4.9 on vector $\mathbf{s}=\mathbf{y}^{\prime}/2$, there exists a vector $\mathbf{t}\in\mathbb{R}^{V}_{\geq 0}$ with $\mathbf{t}\leq\mathrm{deg}_{\partial\mathcal{P}_{i}}/2$ and a flow routing demand $\mathbf{y}^{\prime}/2-\mathbf{t}$ with congestion $\beta$. Scaling this flow by factor $2$, we obtain a flow $f_{2}$ routing demand $\mathbf{y}^{\prime}-2\mathbf{t}$ with congestion $2\beta$.
The final flow is the sum of flows $f_{1}$ and $f_{2}$, which routes demand $(\mathbf{x}^{\prime}-\mathbf{y}^{\prime})+(\mathbf{y}^{\prime}-2\mathbf{t})$ and has congestion $(2L-2(i+1))\beta+2\beta=(2L-2i)\beta$. We set $\mathbf{y}=\mathbf{x}-\mathbf{x}^{\prime}+2\mathbf{t}$ so that the demand routed is exactly $\mathbf{x}-\mathbf{y}$. Note that $\mathbf{y}\geq 0$ since $\mathbf{x}^{\prime}\leq\mathbf{x}$ holds by Claim 4.3.
To complete the induction, it remains to establish $\mathbf{y}\leq 2\deg_{\partial\mathcal{P}_{i}}$. Since $\mathbf{y}=\mathbf{x}-\mathbf{x}^{\prime}+2\mathbf{t}$ and $\mathbf{t}\leq\mathrm{deg}_{\partial\mathcal{P}_{i}}/2$, it suffices to show that $\mathbf{x}-\mathbf{x}^{\prime}\leq\deg_{\partial\mathcal{P}_{i}}$. If $\deg_{\partial\mathcal{R}_{\geq i}}(v)=0$, then $\mathbf{x}(v)=0$ and $\mathbf{x}^{\prime}(v)=0$, so $(\mathbf{x}-\mathbf{x}^{\prime})(v)=0\leq\deg_{\partial\mathcal{P}_{i}}(v)$. Otherwise, if $\deg_{\partial\mathcal{R}_{\geq i}}(v)>0$, then
$(\mathbf{x}-\mathbf{x}^{\prime})(v)$ $=\bigg{(}\frac{\deg_{\partial\mathcal{R}_{\geq i}}(v)-\deg_{\partial\mathcal{R}_{\geq i+1}}(v)}{\deg_{\partial\mathcal{R}_{\geq i}}(v)}\bigg{)}\mathbf{x}(v)$
$\leq\frac{\deg_{\partial\mathcal{P}_{i}}(v)}{\deg_{\partial\mathcal{R}_{\geq i}}(v)}\mathbf{x}(v)$
$\leq\deg_{\partial\mathcal{P}_{i}}(v),$
where the first inequality holds by Claim 4.4. This completes the induction and hence the proof. ∎
The next subclaim almost proves the desired statement, except that $\mathbf{x}$ and $\mathbf{y}$ are restricted to be nonnegative.
Subclaim 4.11.
@@
For any $i\in[L-1]$, consider any vector $\mathbf{x}\in\mathbb{R}^{V}_{\geq 0}$ with $\mathbf{x}\leq\deg_{\partial\mathcal{R}_{\geq i}}$. There exists a vector $\mathbf{y}\in\mathbb{R}^{V}_{\geq 0}$ such that
@@
    $\mathbf{y}\leq 2\deg_{\partial\mathcal{P}_{i}}+2L\beta\deg_{\partial\mathcal{P}_{i+1}}$.
@@
Apply Subclaim 4.10 on vector $\mathbf{x}$, obtaining a vector $\mathbf{y}\in\mathbb{R}^{V}_{\geq 0}$ which we relabel as $\mathbf{y}^{\prime}$, and a flow $f$ routing demand $\mathbf{x}-\mathbf{y}^{\prime}$ with congestion $(2L-2i)\beta\leq 2L\beta$. Take a path decomposition of flow $f$ where each vertex is the start of $\mathbf{x}(v)$ total capacity of (potentially empty) paths and the end of $\mathbf{y}^{\prime}(v)$ total capacity of (potentially empty) paths. For each path starting at a vertex $v$ in some cluster $C\in\mathcal{P}_{i+1}$, perform the following operation. If the path contains an edge $(u,v)$ in $\partial C$ with $u\in C$, then replace the path with its prefix ending at $u$; otherwise, do nothing to the path. Note that the modified path ends in the same cluster $C\in\mathcal{P}_{i+1}$ as its starting point. These modified paths form a new flow $f^{\prime}$, which also has congestion $2L\beta$.
@@
Define vector $\mathbf{y}\in\mathbb{R}^{V}_{\geq 0}$ such that each vertex $u\in V$ is the endpoint of $\mathbf{y}(u)$ total capacity of paths in the new flow $f^{\prime}$. In other words, the flow $f^{\prime}$ routes demand $\mathbf{x}-\mathbf{y}$. We have shown that $\mathbf{y}\leq\mathbf{y}^{\prime}+2L\beta\deg_{\partial\mathcal{P}_{i+1}}$, and combined with the guarantee $\mathbf{y}^{\prime}\leq 2\deg_{\partial\mathcal{P}_{i}}$ from Subclaim 4.10, we conclude that $\mathbf{y}\leq 2\deg_{\partial\mathcal{P}_{i}}+2L\beta\deg_{\partial\mathcal{P}_{i+1}}$. Finally, recall that each path of $f^{\prime}$ starts and ends in the same cluster of $\mathcal{P}_{i+1}$, so $(\mathbf{x}-\mathbf{y})(C)=0$ for all clusters $C\in\mathcal{P}_{i+1}$. ∎
Finally, we prove Claim 4.8 using Subclaim 4.11. Given vector $\mathbf{x}\in\mathbb{R}^{V}$ with $|\mathbf{x}|\leq\deg_{\partial\mathcal{R}_{\geq i}}$, let $\mathbf{x}^{+},\mathbf{x}^{-}\in\mathbb{R}^{V}_{\geq 0}$ be the positive and negative parts of $\mathbf{x}$, so that $\mathbf{x}=\mathbf{x}^{+}-\mathbf{x}^{-}$. Apply Subclaim 4.11 to $\mathbf{x}^{+}$ and $\mathbf{x}^{-}$ separately to obtain $\mathbf{y}^{+}$ and $\mathbf{y}^{-}$, respectively, and set $\mathbf{y}=\mathbf{y}^{+}-\mathbf{y}^{-}$. The three properties are immediately satisfied; note that the congestion is now $4L\beta$ since we take the difference of the two flows routing demands $\mathbf{x}^{+}-\mathbf{y}^{+}$ and $\mathbf{x}^{-}-\mathbf{y}^{-}$. This concludes the proof of Claim 4.8. ∎
@@
Since $|\mathbf{s}-\mathbf{t}|\leq|\mathbf{s}|+|\mathbf{t}|\leq\deg_{\partial\mathcal{R}_{\geq i}}+\deg_{\partial\mathcal{R}_{\geq i+1}}\leq 2\deg_{\partial\mathcal{R}_{\geq i}}$, we apply Claim 4.8 with $\mathbf{x}=\frac{1}{2}(\mathbf{s}-\mathbf{t})$, obtaining a vector $\mathbf{y}\in\mathbb{R}^{V}_{\geq 0}$ with $|\mathbf{y}|\leq 2\deg_{\partial\mathcal{P}_{i}}+2L\beta\deg_{\partial\mathcal{P}_{i+1}}$ and $(\frac{1}{2}(\mathbf{s}-\mathbf{t})-\mathbf{y})(C)=0$ for all clusters $C\in\mathcal{P}_{i+1}$. Let $f_{1}$ be the flow scaled by factor $2$, which routes demand $\mathbf{s}-\mathbf{t}-2\mathbf{y}$ with congestion $8L\beta$.
Consider a cluster $C\in\mathcal{P}_{i+1}$. Since $(\frac{1}{2}(\mathbf{s}-\mathbf{t})-\mathbf{y})(C)=0$ and $(\mathbf{s}-\mathbf{t})(C)=0$, we have $\mathbf{y}(C)=0$ as well. Moreover, for all vertices $v\in C$, we have
$|\mathbf{y}(v)|$ $\leq 2\deg_{\partial\mathcal{P}_{i}}(v)+2L\beta\deg_{\partial C}(v)\leq 4L\beta\deg_{\partial\mathcal{P}_{i}\cup\partial C}(v).$
We conclude that the scaled-down and restricted vector $\frac{1}{4L\beta}\mathbf{y}|_{C}$ is a demand satisfying $\big{|}\frac{1}{4L\beta}\mathbf{y}|_{C}\big{|}\leq\mathrm{deg}_{\partial\mathcal{P}_{i}\cup\partial C}|_{C}$. By assumption (2) of Lemma 4.2, the collection of vertex weightings $\{\mathrm{deg}_{\partial\mathcal{P}_{i}\cup\partial C}|_{C}\in\mathbb{R}^{V}_{\geq 0}:C\in\mathcal{P}_{i+1}\}$ mixes simultaneously in $G$ with congestion $\alpha$, so there is a flow in $G$ routing demand $\sum_{C\in\mathcal{P}_{i+1}}\frac{1}{4L\beta}\mathbf{y}|_{C}=\frac{1}{4L\beta}\mathbf{y}$ with congestion $\alpha$. Scaling this flow by factor $8L\beta$, we obtain a flow $f_{2}$ routing demand $2\mathbf{y}$ with congestion $8L\alpha\beta$.
The final flow $f$ is $f_{1}+f_{2}$, which routes demand $(\mathbf{s}-\mathbf{t}-2\mathbf{y})+2\mathbf{y}=\mathbf{s}-\mathbf{t}$ and has congestion $8L\beta+8L\alpha\beta\leq 16L\alpha\beta$, concluding the proof of Lemma 4.6. ∎
@@ 5 Partitioning Algorithm
The guarantee $\delta\mathcal{P}_{i+1}\leq\frac{1}{2}\delta\mathcal{P}_{i}$ ensures that for $L=O(\log(nW))$, we must have $\delta\mathcal{P}_{L}<1$. Since all edge capacities are assumed to be at least $1$, we conclude that $\delta\mathcal{P}_{L}=0$ and $\mathcal{P}_{L}=\{V\}$, fulfilling property (1) of Theorem 4.1. By Theorem 4.1, we obtain a congestion-approximator with quality $16L^{2}\alpha\beta=O(\log^{10}(nW))$. ∎
@@
At a high level, our algorithm proceeds similarly to expander decomposition algorithms like [SW19], where expanders are defined with respect to the vertex weighting $\deg_{\partial\mathcal{P}_{L}}$ of partition $\mathcal{P}_{L}$. We iteratively decompose the graph using the cut-matching game of [KRV09a]: for each cluster of the decomposition, we either compute a low-conductance cut or certify that the current cluster is mixing (or in [SW19] terms, a nearly expander). The matching step of the cut-matching game requires a call to approximate min-cut/max-flow, but recall that the partitions $\mathcal{P}_{1},\ldots,\mathcal{P}_{L}$ only form a pseudo-congestion-approximator. Luckily, we can modify the graph in a way that the pseudo-congestion-approximator can be adapted to an actual congestion-approximator for the new graph. We then black-box a cut/flow algorithm on this new graph, and then modify the cut and flow to fit the old graph in a way that suffices for the cut-matching game.
@@
    In Section 5.2, we cite the (approximate) fair cut/flow algorithm of [LNPS23], which computes a cut/flow pair with desirable properties given a congestion-approximator.
@@
    In Section 5.3, we introduce the cut-matching game as well as a trimming procedure similar to [SW19]. Both the cut-matching game and the trimming step use the fair cut/flow algorithm (Section 5.2) on the modified graphs for which we have a congestion-approximator (Section 5.1).
@@ 5.1 Congestion-approximator
For any demand $\mathbf{b}\in\mathbb{R}^{V}$ satisfying $|\mathbf{b}(C)|\leq\delta C$ for all $C\in\mathcal{C}$, there exists a demand $\mathbf{b}^{\prime}\in\mathbb{R}^{V}$ satisfying $|\mathbf{b}^{\prime}|\leq\deg_{\partial\mathcal{P}_{L}}$ and a flow in $G$ routing demand $\mathbf{b}-\mathbf{b}^{\prime}$ with congestion $\kappa$, where we define $\kappa=16L^{2}\alpha\beta$.
Suppose we start with the entire graph $G$, and then add a vertex $x$ connected to each vertex $v\in V$ with an edge of capacity $\deg_{\partial\mathcal{P}_{L}}(v)$. In the setting of Fact 5.4, suppose we wish to route the demand $\mathbf{b}$. We start with the flow routing demand $\mathbf{b}-\mathbf{b}^{\prime}$ as promised by Fact 5.4. To route the remaining demand $\mathbf{b}^{\prime}$, we simply use the new edges incident to $x$: for each vertex $v\in V$, send $|\mathbf{b}^{\prime}(v)|\leq\deg_{\partial\mathcal{P}_{L}}$ flow along the edge $(x,v)$ in the proper direction, which is a flow with congestion $1$. Overall, we obtain a flow routing demand $\mathbf{b}$ with congestion $\kappa+1$, and with some more work, we can show that $\mathcal{C}$ is a congestion approximator of the new graph.
@@
Let $\mathcal{C}|_{A}$ be the collection $\{C\cap A:C\in\mathcal{C}\}$. Then, $\mathcal{C}|_{A}\cup\{\{x\},\{s\},\{t\}\}$ is a congestion-approximator of $H$ with quality $O(\beta\gamma^{-1}\kappa)$.
For the rest of Section 5.1, we prove Lemma 5.5. Consider a demand $\mathbf{b}\in\mathbb{R}^{A\cup\{x,s,t\}}$ satisfying $|\mathbf{b}(C\cap A)|\leq\delta_{H}(C\cap A)$ for all $C\in\mathcal{C}$ as well as $|\mathbf{b}(v)|\leq\delta_{H}\{v\}$ for $v\in\{x,s,t\}$. We want to establish a flow on $H$ routing demand $\mathbf{b}$ with congestion $O(\beta\gamma^{-1}\kappa)$.
@@
so we can apply Fact 5.4 on demand $\widetilde{\mathbf{b}}^{\prime}/(1+2\beta+2\gamma)$ to obtain a demand $\mathbf{b}^{\prime}\in\mathbb{R}^{V}$ satisfying $|\mathbf{b}^{\prime}|\leq\deg_{\partial_{G}\mathcal{P}_{L}}$ and a flow on $G$ routing demand $\widetilde{\mathbf{b}}^{\prime}/(1+2\beta+2\gamma)-\mathbf{b}^{\prime}$ with congestion $\kappa$. Scaling this flow by factor $(1+2\beta+2\gamma)$, we obtain a flow $f^{\prime}$ on $G$ routing demand $\widetilde{\mathbf{b}}^{\prime}-(1+2\beta+2\gamma)\mathbf{b}^{\prime}$ with congestion $(1+2\beta+2\gamma)\kappa$.
@@
Finally, since demand $(1+2\beta+2\gamma)\mathbf{b}^{\prime}$ satisfies $|(1+2\beta+2\gamma)\mathbf{b}^{\prime}|\leq(1+2\beta+2\gamma)\deg_{\partial_{G}\mathcal{P}_{L}}$, we can directly route the demand along the edges $(v,x)$ of capacity $\gamma\deg_{\partial_{G}\mathcal{P}_{L}\cup\partial_{G}A}(v)$, which is a routing with congestion $\gamma^{-1}(1+2\beta+2\gamma)$.
Adding up all three routings, we have routed the initial demand $\mathbf{b}$ with congestion $1+\gamma^{-1}(1+2\beta+2\gamma)\kappa+\gamma^{-1}(1+2\beta+2\gamma)=O(\beta\gamma^{-1}\kappa)$, concluding the proof.